Reports 1-1 of 1 Clear search Modify search
terrence.tsang - 12:00 Friday 19 July 2019 (9579) Print this report
SRM QPD centering

With Valera, Kokeyama-san, Fabian,

We noticed that the SRM QPD was not in a good position with cavity/IFO is aligned. This disabled and type of controls that is related to the OL. Therefore, we decided re-center the QPDs today. We aligned the SRM optic. After that, we removed the weights (masses, flanges) that were lying on the SRM frame. After that, we checked the status of SRM again and confirmed that there was no drastic change and that the optic was still aligned. Then, I proceeded to to center the SRM QPDs and saved the new values as the GOOD_OPLEV positions.

When the SRM is not aligned, we noticed that the QPD counts were actually off range. Therefore, we decided to steer the optic with IM pico (for pitch) and F0 yaw stepper (for yaw) to do a coarse alignment according to the new setpoints when the optic is not aligned. This makes the QPDs usable before the optic is aligned.  

Search Help